OpenStudy (anonymous):

*MEDAL* A circle has a diameter of 2.5 ft. What is the area of the circle, rounded to the nearest tenth? Use 3.14 for π.

OpenStudy (igreen):

Formula: \(A_{circle} = \pi r^2\) \(A_{circle} = (3.14) (1.25)^2\) Simplify

OpenStudy (igreen):

The radius is 1.25, because the radius is half the diameter, and since the diameter is 2.5, the radius will be 2.5 / 2 or 1.25.

OpenStudy (igreen):

Multiply 3.14 * 1.25 * 1.25 @Nina_Pooh